For the following unknown, annotate the corresponding spetras and determine the molecular formula and lewis structure.
Anknown IR/HNMR/CNMR MW=122 : 124 AMU 1:1 D 3000 1000 00 RYENUEERI
D, 6 M, 1 10 7 4 1 HP-03-172 Ppm N 10
120 60 100 80 40 20 200 180 160 140 ppm CDS-00-26O
Answer #1

Based on the Spectra, the compound is

2-bromopropane Br 14.29

1H-NMR of 2-bromopropane

1.71, d. 6H 4.29, m, 1H PPM

2-bromopropane have two chemically different proton, first methyl proton give signal towards upfield around 1.4ppm, all six hydrogen are chemically equivalent, splitting pattern is doublet (2nI+1 = 2*1*1/2+1 =2) and another is methine proton which give signal towards downfield around 4.29 ppm, splitting pattern is multiplet (2nI+1 = 2*6*1/2+1 =7).

13C-NMR of 2-bromopropane

29.0 29.0,20 44.5, 1C 45 46 35 36 25 PPM 20 15 16

2-bromopropane has two chemically different carbon, one methyl carbon which shows signal around 30ppm and methine carbon shows peaks at 45ppm.

IR of 2-bromopropane:

C-Br str C-C str sp? C-H str (600

IR spectrum of 2-bromopropane gives three stretching frequency

1. 2900 cm-1 for sp3 C-H stretching

2. 1200 cm-1 for C-C stretching

3. 650 cm-1 for C-Br stretching

Calculated Molecular weight of 2-bromopropane (using 79Br) = 122 g/mol

Given MW = 122 : 124 AMU 1:1

m/z: 121.97 (100,0%), 123.97 (97.3%)

It shows that presence of isotopes with 1:1 ratio, Br isotopes has 100:99.9 ratio, so Br attach to the methine carbon.

Molecular formula of 2-bromopropane = C3H7Br

Lewis structure of C3H7Br:

- H. .c .H H.C : I. . I. Lewis Structure of 2-bromopropane

H. Br: H. Lewis Structure of 2-bromopropane